OpenStudy (anonymous):

Assume f(x) and g(x) are differentiable on [1,4]. If f(1)=2, g(1)=-3, f(4)=5, g(4)=0, and Integral[g(x)f'(x),x,1,4]=1, find Integral[f(x)g'(x),x,1,4] A.-7 B. -1 C. 5 I got C, can anyone check my work?

OpenStudy (amistre64):

cant really make out the querstion

OpenStudy (anonymous):

\[\int\limits_{1}^{4}g(x)f'(x) =1\] \[\int\limits_{1}^{4}f(x)g'(x) =?\]

OpenStudy (anonymous):

given that f(1)=2, g(1)=-3, f(4)=5, g(4)=0

OpenStudy (anonymous):

integrate by parts

OpenStudy (anonymous):

Int = [f(x)g(x)] from 4 to 1 - f'(x)g(x)dx from 1 to 4

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

thats f(4)g(4) - f(1)g(1) - 1

OpenStudy (anonymous):

get it>?????

OpenStudy (anonymous):

6-1 = 5

OpenStudy (anonymous):

okay, so you got 5 as well?

OpenStudy (anonymous):

thats the answer mate

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no prob